OpenStudy (amazingmeisnotonfire):

I will give medal if can help! Find the number of real solutions for the equation. x^2+8=0

OpenStudy (amazingmeisnotonfire):

The formula you use is b^2-4ac, but there's no c, so I don't know what to do

OpenStudy (anonymous):

why not are there ,,c" ? what is there 8 than ?

OpenStudy (amazingmeisnotonfire):

@jhonyy9 b, c whatever! I'm just frustrated b/c I can't figure it out and I've been working on it for ages

OpenStudy (anonymous):

ax^2 +bx +c = 0 this is the formula of quadratic - yes so than you can seeing now what is a and what b and what c - yes ?

OpenStudy (amazingmeisnotonfire):

wut

OpenStudy (anonymous):

so there is b=0 bc bx dont exist - yes ?

OpenStudy (amazingmeisnotonfire):

Yeah I guess

OpenStudy (anonymous):

or you can rewrite this equation same with x^2 +0x +8 = 0 and solve it now in this way - ok. ?

OpenStudy (amazingmeisnotonfire):

So, 0 solutions?

OpenStudy (anonymous):

why zero ?

OpenStudy (abdullahm):

0 real solutions, yes.

OpenStudy (anonymous):

zero real solutions you think ? so than how need being the discriminant ?

OpenStudy (amazingmeisnotonfire):

I only needed to figure out how many solutions there are, so my question is answered. Thanks!

OpenStudy (anonymous):

so than you like check it for zero real solutions the discriminant need being samler than zero x^2 +0x +8 = 0 D = b^2 -4ac D = 0^2 -4*1*8 = -32 so what mean that has zero real solutions hope this is now understandably right easy sure

OpenStudy (abdullahm):

It was my pleasure!